If f(1)=8 and f(n)=1/2f(n-1)+2 then which of the following is the value of f(3)?
Answer by amarjeeth123(569)

f(1)=8
f(2)=1/2f(1)+2=8/2+2=4+2=6
f(3)=1/2f(2)+2=6/2+2=3+2=5
The value of f(3) is 5.